OpenStudy (cwrw238):

I'm not sure how to solve this equation y' (2x - 2y - 4) = 2x + 7y + 5

OpenStudy (cwrw238):

we are studying linear differential equations at the moment

OpenStudy (swissgirl):

Umm what methods do you know as of yet

OpenStudy (swissgirl):

Are you familiar with the method Exact equations?

OpenStudy (cwrw238):

we've been doing integrating factors and homogeneous equations, this question was the last one the set homework.

OpenStudy (cwrw238):

exact equations? no

OpenStudy (swissgirl):

How about substitutions?

OpenStudy (cwrw238):

ah - it was mentioned in our last lecture - do you think it would help here?

OpenStudy (swissgirl):

Ya I do think so actually

OpenStudy (swissgirl):

Wait let me see one sec

OpenStudy (cwrw238):

i'll havev to check in my written notes.

OpenStudy (cwrw238):

i must admit i'm struggling a bit with this topic

OpenStudy (swissgirl):

You can definitely use substitution just gotta figure it out

OpenStudy (swissgirl):

I am having a hard time finding a substitution

OpenStudy (phi):

you can write this equation as (2x - 2y - 4) dy = (2x + 7y + 5) dx have you seen the technique where you shift the origin so that the coefficients are zero: solve 2x - 2y - 4 =0 2x + 7y =0 to find (1,-1) let u= x -1 and v = y+1 to get (2u - 2v) dv = (2u + 7v) du this can be solved using another substitution: v = w u , dv = w du + u dw

OpenStudy (phi):

another method is use the substitution u = 2x -2y -4 du = 2 dx -2 dy v = 2x + 7y + 5 and dv = 2 dx + 7 dy solve for dx and dy in terms of du and dv this should get you to something you can solve